OpenStudy (anonymous):

function f(x) = -2x2, if the domain is {-3, 0, 3}, find the range

jimthompson5910 (jim_thompson5910):

plug in each value in the domain to get corresponding values in the range

jimthompson5910 (jim_thompson5910):

so for instance, if x = -3, then.. f(x) = -2x^2 f(-3) = -2(-3)^2 f(-3) = -2(9) f(-3) = -18

jimthompson5910 (jim_thompson5910):

So x = -3 in the domain corresponds to y = -18 in the range

jimthompson5910 (jim_thompson5910):

Do this for every value in the domain to build the complete range for this function (and restrictions)

OpenStudy (anonymous):

Thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome